Manufacturing Overhead Allocation Base and Calculating the Cost of Jobs. Brenner Company expects to incur $3,000,000 in manufacturing overhead costs this year. During the year, it expects to use 50,000 direct labor hours at a cost of $600,000 and 80,000 machine hours.

Required:

  1. Prepare a predetermined overhead rate based on direct labor hours, direct labor cost, and machine hours.

  2. Why might Brenner Company prefer to use machine hours to allocate manufacturing overhead?

  3. Using each of the predetermined overhead rates calculated in part a and the data that follows for job 128, determine the cost of job 128.

Direct materials $6,000
Direct labor $4,000 (200 hours at $15 per hour) + (100 hours at $10 per hour)
Machine time 700 hours
Answer #1

a) Predetermine overhead rate

Base on direct labor hour = 3000000/50000 = 60 per labor hour

Direct labor cost = 3000000/600000 = $5 per dollar of labor cost

Machine hour = 3000000/80000 = 37.50 per machine hour

b) Job Cost

Base on labor hour Base on labor cost Base on machine hour
Direct material 6000 6000 6000
Direct labor 4000 4000 4000
Overhead 300*60 = 18000 4000*5 = 20000 700*37.50 = 26250
Total 28000 30000 36250

c) Brenner use machine hour because most of the companies activity by machining hour so brenner use machine hour for allocation